OpenStudy (anonymous):

Which are measurements of the sides of a right triangle? A. 202, 198, 40 B. 28, 26, 12 C. 17, 14, 6 D. All above?

OpenStudy (nottim):

c2=a2+b2

OpenStudy (nottim):



OpenStudy (anonymous):

Anything where the (longest side)^2 = (one side)^2 + (other side)^2.

OpenStudy (anonymous):

so is it C.?

OpenStudy (anonymous):

Does 17^2 = 14^2 + 6^2?

OpenStudy (anonymous):

dang so which is it?

sam (.sam.):

A

OpenStudy (anonymous):

na that cant be it right

sam (.sam.):

\[\huge \sqrt{198^{2}+40^{2}}=202\] Others are not equal